Complete the following fusion reaction: 2H+3H⟶n+−−−−−−.

Express your answer as a complete nuclear equation.

nothing

SubmitRequest Answer
Answer #1

The complete reaction is:

_{1}^{2} extrm{H} + _{1}^{3} extrm{H} ightarrow _{2}^{4} extrm{He} + _{0}^{1} extrm{n} + 17.6 MeV

Solution:

We need to evaluate the mass decrease in the above reaction.

H(2.014102) HH(3.016050) →Яне(4.002603) n(1.008665)

Total mass on reactant side = 5.030152

Total mass on product side = 5.011268

Mass defect = 5.030152 - 5.011268 = 0.018884 a.m.u. are converted to energy

small Delta E = mc^2

1.66056 x 1027ka 018884amu amu

17.6MeV/nucleus